Hi all, this is my last one.
You buy a baseball card for 15 cents, sell it for 55 cents, buy it back for 40 cents and sell it again for a dollar. How much did you make?
Thanks you
Gina

Found 2 solutions by Theo, josmiceli:
Answer by Theo(13342)

buy it for 15 cents.
sell it for 55.
buy it back for 40 cents.
sell it for a dollar.
your cash flow is:
-15 + 55 - 40 + 100
you spent a total of 55 cents and got back a total of 155 cents.
your profit is 155 - 55 = 100.
you made 40 cents on the first transaction and you made 60 cents on the second transaction which means you made a total of 100.

Answer by josmiceli(19441)

You account balances are:
+15
15 - 15 = 0
+55
55 - 40 = 15
15 + 100 = 115
You started with 15 cents and you ended up
with $1.15, so you made $1.00